OpenStudy (anonymous):

The solutions of : √a - 1/b - 2c + 3d = 1 a + 1/b² - 4c² - 9d² = 3 a√a - 1/b³ - 8c³ + 27d³ = -5 a² + 1/b⁴ -16c⁴ - 81z⁴ = 15 is ...?

OpenStudy (anonymous):

\[\sqrt{a}-\frac{1}{b}-2c+3d=1\]\[a-\frac{1}{b^2}-4c^2-9d^2=3\]\[a\sqrt{a}-\frac{1}{b^3}-8c^3+27d^3=-5\]\[a^2+\frac{1}{b^4}-16c^4-81d^4=15\] ?

OpenStudy (anonymous):

let \[A=\sqrt{a}\]\[B=\frac{1}{b}\]\[C=2c\]\[D=3d\]and see here http://openstudy.com/study#/updates/5025344ce4b0d38989b7378f

OpenStudy (anonymous):

@tanjung

OpenStudy (anonymous):

waw, very spectaculer...thanku very much mukushla

OpenStudy (anonymous):

yw :) this is a nice problem